# Generalized Virial Identities: Radial Constraints for Solitons, Instantons, and Bounces

### v1

@@ Abstract
Topological solitons and tunneling configurations (monopoles, vortices, Skyrmions, sphalerons, and bounces) are characterized by profile functions that encode their physical properties. Derrick’s scaling relation provides a global integral constraint on these profiles, but reveals nothing about the local balance between kinetic and potential contributions in different regions. We derive a continuous family of virial identities for O($n$) symmetric configurations, parameterized by an exponent $\alpha$ that controls the radial weighting: negative $\alpha$ emphasizes the core where topological boundary conditions are imposed, large positive $\alpha$ emphasizes the asymptotic tail, and $\alpha=1$ recovers the classical Derrick relation. The family provides a systematic decomposition of the global constraint into radially-resolved components, with special $\alpha$ values isolating specific mechanisms. For BPS configurations, where the Bogomolny equations imply pointwise equality between kinetic and potential densities, the virial identity is satisfied for all valid $\alpha$. We verify the formalism analytically for the Fubini-Lipatov instanton, BPS monopole, and BPST instanton. Numerical tests on the Coleman bounce and Nielsen-Olesen vortex illustrate how the $\alpha$-dependence of errors distinguishes core from tail inaccuracies: the vortex shows errors growing at negative $\alpha$ (core), while the bounce shows errors growing at positive $\alpha$ (tail). Applications to the electroweak sphaleron, where the Higgs mass explicitly breaks scale invariance, and the hedgehog Skyrmion illustrate the formalism in systems with multiple competing length scales.
@@ 1 Introduction
Our work connects to several recent developments in integral identities for solitons. Manton Manton2009 derived scaling identities beyond Derrick’s theorem by considering more general coordinate transformations, applying infinitesimal diffeomorphisms to the action functional. Our $\alpha$-family overlaps with Manton’s identities at specific parameter values but differs in emphasis: we parameterize the identities continuously and focus on their utility as radial probes. Adam et al. Adam2024 demonstrated that for any nonlinear field theory supporting static solutions with finite energy, an infinite number of integral identities can be derived from field transformations and their Noether currents. They showed that identities generated by coordinate transformations become trivial for BPS solitons, while field-space transformations yield nontrivial relations connected to observables such as the $D$-term and mechanical radius PolyakovSchweitzer2018; Ji2021. Our formalism provides explicit realizations of coordinate-based identities for O($n$) symmetric configurations and confirms the BPS triviality. Gudnason and Speight GudnasonSpeight2017 established a related 3-parameter family of integral identities for spherically symmetric solitons using boundary charge methods; our approach differs in deriving the identities from the radial Hamiltonian structure. Herdeiro et al. HerdeiroOliveiraPomboRadu2021; HerdeiroOliveiraPomboRadu2022 extended virial identities to relativistic gravity, showing how they can be recast as energy-momentum balance conditions for self-gravitating solitons and black holes. The $\alpha$-family also generalizes the classical Pohozaev identities Pohozaev1965 from elliptic PDE theory: Pohozaev derived integral constraints by multiplying the equation $\Delta u=f(u)$ by $x\cdot\nabla u$ and integrating by parts, which for radial solutions corresponds to $\alpha=1$. Our generalization to arbitrary $\alpha$ yields a continuous family of constraints.
@@ 2.1 Radial Hamiltonian
where $\Omega_{n-1}=2\pi^{n/2}/\Gamma(n/2)$ is the area of the unit (n1)(n-1)-sphere and $\mathcal{G}_{\rho}$ is the reduced integrand. The geometric factor $\rho^{n-1}$ from the volume element $d^{n}x=\rho^{n-1}d\rho\,d\Omega_{n-1}$ introduces explicit $\rho$-dependence into $\mathcal{G}_{\rho}$, even when the original Lagrangian density $\mathcal{G}$ has no such dependence. This explicit $\rho$-dependence is the origin of all virial constraints.
@@
which is the Legendre transform of $\mathcal{G}_{\rho}$ with respect to the velocities, i.e., the “radial Hamiltonian” of the one-dimensional system. Computing its derivative along a solution:
@@
The rate of change of the radial Hamiltonian equals (minus) the explicit $\rho$-derivative of the Lagrangian. If $\mathcal{G}_{\rho}$ had no explicit $\rho$-dependence, $\mathcal{C}_{\rho}$ would be conserved, the analogue of energy conservation in time-independent mechanics. The geometric weights break this conservation, and Eq. (5) quantifies the breaking.
@@ 2.6 Derrick’s theorem
By contrast, $\alpha<1$ (and especially negative $\alpha$) shifts the weight toward small $\rho$. The factor $\rho^{\alpha-1}$ in the kinetic integrand diverges as $\rho\to 0$ for $\alpha<1$, amplifying the contribution from the core. If finite-difference errors are concentrated where the field gradients are steepest (typically near the core), these errors are amplified by negative-$\alpha$ weighting. The result is a virial discrepancy that the standard $\alpha=1$ test misses entirely. The $\alpha$-family thus provides independent constraints that probe different radial regions.
@@ 3.1 Fubini-Lipatov instanton
This solution arises in four-dimensional Euclidean scalar field theory with the conformal potential
$V(\phi)=-\frac{\lambda}{4}\phi^{4}.$ (16)
Although this potential is unbounded from below (raising questions of stability that we set aside here), it admits finite-action O(4) symmetric solutions relevant to conformal anomalies and certain cosmological scenarios. The exact instanton solution, discovered independently by Fubini and Lipatov Rajaraman1982, takes the form
@@ 3.2 Vacuum decay
The bounce configuration interpolates between the false vacuum at spatial infinity and a region near the true vacuum at the origin, representing the nucleation of a critical bubble. We consider the quartic potential with explicit symmetry breaking,
$V(\phi)=\frac{\lambda}{8}(\phi^{2}-a^{2})^{2}-\frac{\epsilon}{2}(\phi+a),$ (26)
which for small $\epsilon>0$ has a metastable false vacuum near $\phi_{\rm fv}\approx-a$ and a deeper true vacuum near $\phi_{\rm tv}\approx+a$. The bounce solution satisfies the Euclidean equation of motion with boundary conditions $\dot{\phi}(0)=0$ (regularity at the origin) and $\phi(\infty)=\phi_{\rm fv}$ (approach to the false vacuum).
@@
coincides with the “mechanical energy” in Coleman’s analogy  Coleman1977. The evolution equation (5) at $\alpha=-3$ becomes
@@ 4.1 BPS monopole
The ’t Hooft-Polyakov monopole tHooft1974; Polyakov1974 offers a stringent test of the virial formalism in gauge theories. This solution arises in SU(2) Yang-Mills-Higgs theory, where a triplet Higgs field breaks the gauge symmetry to U(1). In the BPS limit Bogomolny1976, where the Higgs self-coupling vanishes, the monopole saturates a topological energy bound and admits an exact analytical solution. The reduced energy functional takes the form
$E=\frac{4\pi v}{g}\int_{0}^{\infty}\left[(\dot{K})^{2}+\frac{(K^{2}-1)^{2}}{\rho^{2}}+(\rho\dot{H})^{2}+K^{2}H^{2}\right]d\rho,$ (36)
where $\rho$ is the dimensionless radial coordinate (in units of $(gv)^{-1}$), $K(\rho)$ is the gauge profile function, and $H(\rho)$ is the Higgs profile.
The four terms in (36) have distinct origins and geometric weights. The gauge kinetic term $(\dot{K})^{2}$ represents the energy cost of radial variation in the gauge field, carrying unit weight $A_{K}=1$. The term $(K^{2}-1)^{2}/\rho^{2}$ arises from the non-Abelian magnetic field; the factor $1/\rho^{2}$ acts as a centrifugal barrier that diverges as $\rho\to 0$, enforcing the boundary condition $K(0)=1$ and preventing the gauge field from collapsing to its vacuum value $K=0$ prematurely. The Higgs kinetic term $(\rho\dot{H})^{2}$ has weight $A_{H}=\rho^{2}$, reflecting the angular components of the covariant derivative. The coupling term $K^{2}H^{2}$ represents gauge-Higgs interaction energy.
@@ 4.2 BPST instanton
The BPST instanton BPST1975 in four-dimensional pure Yang-Mills theory offers another example where conformal invariance renders the standard Derrick test trivial, while the $\alpha\neq 1$ relations remain nontrivial. The O(4) symmetric ansatz reduces the action to
$S[w]=\frac{24\pi^{2}}{g^{2}}\int_{0}^{\infty}\left[\frac{1}{2}\rho\dot{w}^{2}+\frac{2}{\rho}w^{2}(1-w)^{2}\right]d\rho,$ (42)
where $w(\rho)$ is a profile function satisfying $w(0)=0$ and $w(\infty)=1$.
The two terms display the characteristic kinetic-potential competition. The kinetic term $\frac{1}{2}\rho\dot{w}^{2}$ represents the action density from radial variation, with weight $A=\frac{1}{2}\rho$ from the four-dimensional measure. The potential term $\frac{2}{\rho}w^{2}(1-w)^{2}$ arises from gauge field self-interaction; the $1/\rho$ weight acts as a centrifugal-like barrier that concentrates action density near the instanton center. The competition between the outward “pressure” from the kinetic term and the inward “confinement” from the potential term determines the instanton profile.

### v2

@@ Abstract
We derive a continuous family of virial identities for O($n$) symmetric configurations, parameterized by an exponent $\alpha$ that controls the radial weighting. The family provides a systematic decomposition of the global constraint into radially-resolved components, with special $\alpha$ values isolating specific mechanisms. For BPS configurations, where the Bogomolny equations imply pointwise equality between kinetic and potential densities, the virial identity is satisfied for all valid $\alpha$. We verify the formalism analytically for the Fubini-Lipatov instanton, BPS monopole, and BPST instanton. Numerical tests on the Coleman bounce and Nielsen-Olesen vortex illustrate how the $\alpha$-dependence of errors distinguishes core from tail inaccuracies: the vortex shows errors growing at negative $\alpha$ (core), while the bounce shows errors growing at positive $\alpha$ (tail). Applications to the electroweak sphaleron, where the Higgs mass explicitly breaks scale invariance, and the hedgehog Skyrmion illustrate the formalism in systems with multiple competing length scales.
@@ 1 Introduction
Our work connects to several recent developments in integral identities for solitons. Manton Manton2009 derived scaling identities beyond Derrick’s theorem by considering more general coordinate transformations, applying infinitesimal diffeomorphisms to the action functional. Our $\alpha$-family overlaps with Manton’s identities at specific parameter values but differs in emphasis: we parameterize the identities continuously and focus on their utility as radial probes. Adam et al. Adam2024 demonstrated that for any nonlinear field theory supporting static solutions with finite energy, an infinite number of integral identities can be derived from field transformations and their Noether currents. They showed that identities generated by coordinate transformations become trivial for BPS solitons, while field-space transformations yield nontrivial relations connected to observables such as the $D$-term and mechanical radius PolyakovSchweitzer2018; Ji2021. Our formalism provides explicit realizations of coordinate-based identities for O($n$) symmetric configurations and confirms the BPS triviality. Gudnason, Gao, and Yang GudnasonGaoYang2017 established a related 3-parameter family of integral identities for spherically symmetric solitons using boundary charge methods; our approach differs in deriving the identities from the radial functional structure. Herdeiro et al. HerdeiroOliveiraPomboRadu2021; HerdeiroOliveiraPomboRadu2022 extended virial identities to relativistic gravity, showing how they can be recast as energy-momentum balance conditions for self-gravitating solitons and black holes. The $\alpha$-family also generalizes the classical Pohozaev identities Pohozaev1965 from elliptic PDE theory: Pohozaev derived integral constraints by multiplying the equation $\Delta u=f(u)$ by $x\cdot\nabla u$ and integrating by parts, which for radial solutions corresponds to $\alpha=1$. Our generalization to arbitrary $\alpha$ yields a continuous family of constraints.
@@ 2.1 Radial Functional Density
where $\Omega_{n-1}=2\pi^{n/2}/\Gamma(n/2)$ is the area of the unit (n1)(n-1)-sphere and $\mathcal{G}_{\rho}$ is the reduced integrand. The geometric factor $\rho^{n-1}$ from the volume element $d^{n}x=\rho^{n-1}d\rho\,d\Omega_{n-1}$ introduces explicit $\rho$-dependence into $\mathcal{G}_{\rho}$, even when the original functional density $\mathcal{G}$ has no such dependence. This explicit $\rho$-dependence is the origin of all virial constraints.
@@
which is the Legendre transform of $\mathcal{G}_{\rho}$ with respect to the velocities. Computing its derivative along a solution:
@@
The rate of change of the Legendre transform of the functional $\mathcal{C_{\rho}}$ equals (minus) the explicit $\rho$-derivative of the functional. If $\mathcal{G}_{\rho}$ had no explicit $\rho$-dependence, $\mathcal{C}_{\rho}$ would be conserved, the analogue of energy conservation in time-independent mechanics. The geometric weights break this conservation, and Eq. (5) quantifies the breaking.
@@ 2.3 Kinetic-potential decomposition
The procedure for applying this formalism to a specific system is: (i) write the full Lagrangian or energy functional of the field theory; (ii) specify the O($n$) symmetric ansatz that reduces the system to a radial problem; (iii) perform the angular integration to obtain the reduced functional $\mathcal{G}_{\rho}$; (iv) compute the auxiliary quantity $\mathcal{C}_{\rho}$ from (3); (v) apply the fundamental identity (5) and integrate with weight $\rho^{\alpha}$ to obtain the virial constraints. In subsequent sections, we carry out this procedure explicitly for scalar instantons, gauge field configurations, and chiral solitons.
@@ 2.6 Derrick’s theorem
By contrast, $\alpha<1$ (and especially negative $\alpha$) shifts the weight toward small $\rho$. The factor $\rho^{\alpha-1}$ in the kinetic integrand diverges as $\rho\to 0$ for $\alpha<1$, amplifying the contribution from the core. If finite-difference errors are concentrated where the field gradients are steepest (typically near the core), these errors are amplified by negative-$\alpha$ weighting. The result is a virial discrepancy that the standard $\alpha=1$ test. The $\alpha$-family thus provides independent constraints that probe different radial regions.
@@ 3.1 Fubini-Lipatov instanton
This solution arises in four-dimensional Euclidean scalar field theory with action
$S_{E}[\phi]=\int d^{4}x\left[\frac{1}{2}(\partial_{\mu}\phi)^{2}+V(\phi)\right],\qquad V(\phi)=-\frac{\lambda}{4}\phi^{4}.$ (16)
Although this potential is unbounded from below (raising questions of stability that we set aside here), it admits finite-action O(4) symmetric solutions relevant to conformal anomalies and certain cosmological scenarios.
For O(4) symmetric configurations $\phi=\phi(\rho)$ with $\rho=|x|$, the action reduces to
$S_{E}=2\pi^{2}\int_{0}^{\infty}\mathcal{G}_{\rho}\,d\rho,\qquad\mathcal{G}_{\rho}=\rho^{3}\left[\frac{1}{2}\dot{\phi}^{2}+V(\phi)\right],$ (17)
where $\dot{\phi}=d\phi/d\rho$. The auxiliary quantity (3) becomes
$\mathcal{C}_{\rho}=\rho^{3}\left[\frac{1}{2}\dot{\phi}^{2}-V(\phi)\right].$ (18)
The exact instanton solution, discovered independently by Fubini and Lipatov Rajaraman1982, takes the form
@@
The virial identity requires $\rho^{\alpha}\mathcal{C}_{\rho}]_{0}^{\infty}=0$. From the solution (19):
At the origin ($\rho\to 0$): The field approaches a constant $\phi(0)=\phi_{0}/R$, and the derivative vanishes as
$\dot{\phi}=-\frac{2\phi_{0}R\rho}{(\rho^{2}+R^{2})^{2}}\sim-\frac{2\phi_{0}\rho}{R^{3}}\quad\Rightarrow\quad\dot{\phi}^{2}\sim\rho^{2}.$ (28)
The potential $V(\phi)\to V(\phi_{0}/R)=-\lambda\phi_{0}^{4}/(4R^{4})$ remains finite. Thus $\mathcal{C}_{\rho}=\rho^{3}[\frac{1}{2}\dot{\phi}^{2}-V]\sim\rho^{3}$, and
$\rho^{\alpha}\mathcal{C}_{\rho}\sim\rho^{\alpha+3}\to 0\quad\text{for }\alpha>-3.$ (29)
At infinity ($\rho\to\infty$): The field decays as $\phi\sim\phi_{0}R/\rho^{2}$, giving
$\dot{\phi}\sim-\frac{2\phi_{0}R}{\rho^{3}},\quad\dot{\phi}^{2}\sim\rho^{-6},\quad V\sim-\frac{\lambda\phi_{0}^{4}R^{4}}{4\rho^{8}}\sim\rho^{-8}.$ (30)
The kinetic term dominates, so $\mathcal{C}_{\rho}\sim\rho^{3}\cdot\rho^{-6}=\rho^{-3}$, and
$\rho^{\alpha}\mathcal{C}_{\rho}\sim\rho^{\alpha-3}\to 0\quad\text{for }\alpha<3.$ (31)
The identities are valid for $3<\alpha<3$. The finite upper bound reflects the power-law decay characteristic of conformally invariant theories.
@@ 3.2 Vacuum decay
The bounce configuration interpolates between the false vacuum at spatial infinity and a region near the true vacuum at the origin, representing the nucleation of a critical bubble. The Euclidean action is
$S_{E}[\phi]=\int d^{4}x\left[\frac{1}{2}(\partial_{\mu}\phi)^{2}+V(\phi)\right],$ (32)
with the quartic potential featuring explicit symmetry breaking:
$V(\phi)=\frac{\lambda}{8}(\phi^{2}-a^{2})^{2}-\frac{\epsilon}{2}(\phi+a).$ (33)
For small $\epsilon>0$, this potential has a metastable false vacuum near $\phi_{\rm fv}\approx-a$ and a deeper true vacuum near $\phi_{\rm tv}\approx+a$.
For O(4) symmetric configurations $\phi=\phi(\rho)$, the reduced functional and auxiliary quantity are
$\mathcal{G}_{\rho}=\rho^{3}\left[\frac{1}{2}\dot{\phi}^{2}+V(\phi)\right],\qquad\mathcal{C}_{\rho}=\rho^{3}\left[\frac{1}{2}\dot{\phi}^{2}-V(\phi)\right],$ (34)
with $V$ shifted so that $V(\phi_{\rm fv})=0$. The bounce satisfies $\dot{\phi}(0)=0$ (regularity at the origin) and $\phi(\infty)=\phi_{\rm fv}$ (approach to the false vacuum).
@@
coincides with the mechanical energy in Coleman’s analogy  Coleman1977. The evolution equation (5) at $\alpha=-3$ becomes
@@
The bounce interpolates between vacua rather than decaying to zero, changing the asymptotic structure.
At the origin ($\rho\to 0$): Regularity requires $\dot{\phi}(0)=0$, so
$\phi(\rho)=\phi_{0}+\tfrac{1}{2}\phi^{\prime\prime}(0)\rho^{2}+O(\rho^{4}),\quad\dot{\phi}\sim\phi^{\prime\prime}(0)\rho,\quad\dot{\phi}^{2}\sim\rho^{2}.$ (44)
With $V(\phi_{0})$ finite, $\mathcal{C}_{\rho}=\rho^{3}[\frac{1}{2}\dot{\phi}^{2}-V]\sim\rho^{3}$, and
$\rho^{\alpha}\mathcal{C}_{\rho}\sim\rho^{\alpha+3}\to 0\quad\text{for }\alpha>-3.$ (45)
At infinity ($\rho\to\infty$): The field approaches the false vacuum exponentially,
$\phi-\phi_{\rm fv}\sim e^{-m\rho},\quad\dot{\phi}\sim-m\,e^{-m\rho},\quad m^{2}=V^{\prime\prime}(\phi_{\rm fv}).$ (46)
Both $\dot{\phi}^{2}$ and $V-V(\phi_{\rm fv})$ decay as $e^{-2m\rho}$, so $\mathcal{C}_{\rho}\sim\rho^{3}e^{-2m\rho}$. Exponential decay defeats any polynomial weight:
$\rho^{\alpha}\mathcal{C}_{\rho}\sim\rho^{\alpha+3}e^{-2m\rho}\to 0\quad\text{for all finite }\alpha.$ (47)
The virial relations are valid for $\alpha>-3$, with no upper bound. The absence of an upper bound reflects the massive nature of the theory near the false vacuum.
@@ 4.1 BPS monopole
The ’t Hooft-Polyakov monopole tHooft1974; Polyakov1974 offers a stringent test of the virial formalism in gauge theories. This solution arises in SU(2) Yang-Mills-Higgs theory with Lagrangian
$\mathcal{L}=-\frac{1}{4}F^{a}_{\mu\nu}F^{a\mu\nu}+\frac{1}{2}(D_{\mu}\Phi^{a})(D^{\mu}\Phi^{a})-\frac{\lambda}{4}(|\Phi|^{2}-v^{2})^{2},$ (48)
where $\Phi^{a}$ is a triplet Higgs field breaking SU(2) to U(1). In the BPS limit Bogomolny1976 $\lambda\to 0$, the monopole saturates a topological energy bound and admits an exact analytical solution.
The spherically symmetric hedgehog ansatz is
$\Phi^{a}=vH(\rho)\frac{x^{a}}{\rho},\qquad A^{a}_{i}=\epsilon_{aij}\frac{x^{j}}{\rho^{2}}[1-K(\rho)],$ (49)
where $\rho$ is the dimensionless radial coordinate (in units of $(gv)^{-1}$). Substituting into the energy functional yields the reduced form
$E=\frac{4\pi v}{g}\int_{0}^{\infty}\mathcal{G}_{\rho}\,d\rho,\qquad\mathcal{G}_{\rho}=(\dot{K})^{2}+\frac{(K^{2}-1)^{2}}{\rho^{2}}+(\rho\dot{H})^{2}+K^{2}H^{2},$ (50)
where $K(\rho)$ is the gauge profile and $H(\rho)$ is the Higgs profile.
The auxiliary quantity (3) for this two-field system is
$\mathcal{C}_{\rho}=\dot{K}\frac{\partial\mathcal{G}_{\rho}}{\partial\dot{K}}+\dot{H}\frac{\partial\mathcal{G}_{\rho}}{\partial\dot{H}}-\mathcal{G}_{\rho}=(\dot{K})^{2}+\rho^{2}(\dot{H})^{2}-\frac{(K^{2}-1)^{2}}{\rho^{2}}-K^{2}H^{2}.$ (51)
The explicit $\rho$-dependence in $\mathcal{G}_{\rho}$ arises from the geometric weights $\rho^{-2}$ on the magnetic term and $\rho^{2}$ on the Higgs kinetic term, reflecting the angular structure of the hedgehog configuration.
@@
For the BPS monopole, the Bogomolny equations imply $\mathcal{C}_{\rho}=0$ pointwise, so the boundary term vanishes identically. However, the valid $\alpha$ range is still determined by integral convergence.
At the origin ($\rho\to 0$): The regular solution behaves as
$K=1-\frac{\rho^{2}}{6}+O(\rho^{4}),\quad H=\frac{\rho}{3}+O(\rho^{3}).$ (57)
Thus $\dot{K}\sim-\rho/3$, $\dot{H}\sim 1/3$, and each term in $\mathcal{G}_{\rho}$ scales as $\rho^{2}$:
$\dot{K}^{2}\sim\rho^{2},\quad\rho^{2}\dot{H}^{2}\sim\rho^{2},\quad\frac{(K^{2}-1)^{2}}{\rho^{2}}\sim\rho^{2},\quad K^{2}H^{2}\sim\rho^{2}.$ (58)
The integral $I_{\dot{K}}^{\alpha}=\int\dot{K}^{2}\rho^{\alpha-1}d\rho$ has integrand $\sim\rho^{\alpha+1}$, converging for $\alpha>-2$.
At infinity ($\rho\to\infty$): The gauge field decays exponentially while the Higgs field approaches its vacuum algebraically:
$K\sim 2\rho\,e^{-\rho},\quad H\to 1-\frac{1}{\rho}.$ (59)
Since $\mathcal{C}_{\rho}=0$ exactly for the BPS solution, the boundary term vanishes identically. The remaining constraint comes from integral convergence at infinity.
Thus the relations hold for $\alpha>-2$. The pointwise BPS condition $\mathcal{C}_{\rho}=0$ ensures the virial identity is satisfied throughout this range.
@@ 4.2 BPST instanton
The BPST instanton BPST1975 in four-dimensional pure Yang-Mills theory offers another example where conformal invariance renders the standard Derrick test trivial, while the $\alpha\neq 1$ relations remain nontrivial. The Euclidean action is
$S_{E}=\frac{1}{2g^{2}}\int d^{4}x\,\text{Tr}(F_{\mu\nu}F^{\mu\nu}).$ (60)
The O(4) symmetric ansatz for SU(2) gauge fields,
$A_{\mu}^{a}=\frac{2}{g}\bar{\eta}^{a}_{\mu\nu}\frac{x^{\nu}}{\rho^{2}}w(\rho),$ (61)
where $\bar{\eta}^{a}_{\mu\nu}$ is the ’t Hooft symbol, reduces the action to
$S_{E}=\frac{24\pi^{2}}{g^{2}}\int_{0}^{\infty}\mathcal{G}_{\rho}\,d\rho,\qquad\mathcal{G}_{\rho}=\frac{1}{2}\rho\dot{w}^{2}+\frac{2}{\rho}w^{2}(1-w)^{2},$ (62)
where the profile function $w(\rho)$ satisfies $w(0)=0$ and $w(\infty)=1$.
The auxiliary quantity is
$\mathcal{C}_{\rho}=\frac{1}{2}\rho\dot{w}^{2}-\frac{2}{\rho}w^{2}(1-w)^{2}.$ (63)
The kinetic term carries weight $\rho$ from the four-dimensional measure, while the potential term carries $\rho^{-1}$, acting as a centrifugal-like barrier that concentrates action density near the instanton center.
@@
From the BPST solution $w=\rho^{2}/(\rho^{2}+\rho_{0}^{2})$:
At the origin ($\rho\to 0$): The profile vanishes as $w\sim\rho^{2}/\rho_{0}^{2}$, giving
$\dot{w}=\frac{2\rho\rho_{0}^{2}}{(\rho^{2}+\rho_{0}^{2})^{2}}\sim\frac{2\rho}{\rho_{0}^{2}},\quad\dot{w}^{2}\sim\rho^{2}.$ (67)
The potential term: $w^{2}(1-w)^{2}\sim\rho^{4}/\rho_{0}^{4}$. Thus
$\mathcal{C}_{\rho}=\tfrac{1}{2}\rho\dot{w}^{2}-\tfrac{2}{\rho}w^{2}(1-w)^{2}\sim\rho^{3}-\rho^{3}\sim\rho^{3},$ (68)
and $\rho^{\alpha}\mathcal{C}_{\rho}\sim\rho^{\alpha+3}\to 0$ for $\alpha>-3$.
At infinity ($\rho\to\infty$): The profile approaches unity as $1-w\sim\rho_{0}^{2}/\rho^{2}$, giving
$\dot{w}\sim\frac{2\rho_{0}^{2}}{\rho^{3}},\quad\dot{w}^{2}\sim\rho^{-6},\quad w^{2}(1-w)^{2}\sim\frac{\rho_{0}^{4}}{\rho^{4}}.$ (69)
Both terms in $\mathcal{C}_{\rho}$ scale as $\rho^{-5}$:
$\rho\dot{w}^{2}\sim\rho^{-5},\quad\frac{w^{2}(1-w)^{2}}{\rho}\sim\rho^{-5},$ (70)
so $\mathcal{C}_{\rho}\sim\rho^{-5}$ and $\rho^{\alpha}\mathcal{C}_{\rho}\sim\rho^{\alpha-5}\to 0$ for $\alpha<5$.
In this case the realtions hold for $3<\alpha<5$. Like the Fubini-Lipatov instanton, the finite upper bound reflects power-law decay in this conformally invariant theory.